# Quantum-enabled optical large-baseline interferometry: applications, protocols and feasibility

### v1

@@ V Technological requirements of OVLBI
    Collect the starlight and time-multiplex them into different bins (see Fig. 4)
@@
    The collected starlight is absorbed into quantum memories. Non-local measurement is performed between quantum memories of the same time bin between the telescope sites. Given that each shared time-bin contains a state of mean photon number $\epsilon$, the logical gate error rate needs to be much smaller than $\epsilon$. Per time bin, we need at least
@@
      One set of logical single-qubit gates to perform the measurements. This one is non-Clifford in general
Below, we outline these core requirements and the current state of the art.
@@ V.1 Phase stabilisation
Maintaining phase coherence across distant telescope sites is critical for interferometry. The entire system must be stabilised to within a small fraction of the optical wavelength. A shared optical phase reference is necessary. Notably, recent twin-field QKD experiments have achieved 98% interference visibility over 200 km at 1550 nm [36], suggesting that phase stabilisation over VLBI-scale baselines is technically feasible.
@@ V.2 Time-bin multiplexing
Starlight must be parsed into discrete time bins, with each bin corresponding to the coherence time of the incoming thermal light. The coherence time of the light is inversely proportional to the bandwidth $\Delta\nu$ [29],
@@
As the bandwidth increases, the coherence time—and therefore the required time-bin size—decreases proportionally. As an example, for light generated by thermal sources with a narrow spectral width ($\Delta\nu\sim 10^{8}$ Hz), the corresponding coherence time is of order $10^{-8}s$, and coherence length is of order 3 m. For ease of comparison, Table. 5 provides a comparison for the frequency vs wavelength bandwidth and their conversion for various wavelengths.
Current state-of-the-art optical switching operates at rates of up to 38 MHz with approximately 80% transmission efficiency [37], corresponding to a minimum time-bin size of roughly $2.6\times 10^{-8}$ seconds. In contrast, at a wavelength of 1550 nm, a spectral bandwidth of 1 nm corresponds to a temporal coherence time of about 8 picoseconds—nearly 3000 times shorter than the achievable switching interval. This mismatch between the coherence times and switching capabilities is a challenge for time-bin multiplexing at optical frequencies.
@@ V.3 Implementing quantum nodes with optical cavities
The many platforms for implementing quantum nodes in physical systems are built around memory qubits, which store the information in long-lived nuclear spin, dark or metastable states, perfectly isolated from the environment. To control and read out that memory, quantum nodes would typically use a direct, controlled coupling to a cavity, or mediate it via a short-lived communication qubit [39]. The cavity is used to boost the intrinsically low efficiency of these processes and enhance the in- and out-coupling from the memory to the fibre.
To this end, the cavity operates in the overcoupled regime, meaning that its coupling to the fiber mode $\kappa_{\text{out}}$ should dominate over other losses $\kappa_{0}$, and support high cooperativity $C=g^{2}/\kappa\gamma\gg 1$, defined by the cavity-qubit coupling $g$, and decoherence rate of the qubit $\gamma$.
@@ V.3.1 Atomic defects, incorporating nitrogen, silicon, and germanium atoms
Atomic defects, incorporating nitrogen, silicon, or germanium atoms next to vacancies in diamond (NV, SiV, and GeV) can implement both the communication and memory qubits, encoding information in the electronic transitions of the defect, nuclear nuclear spins of the nearby atoms—13C for NVs, and 29Si for SiVs—respectively [41, 42]. These two types of defects also offer very different challenges and opportunities: NV exhibits far longer coherence times at room temperature ($\sim 1$ s, several orders of magnitude over SiV), but also is far more sensitive to electrical noise and coupling to phonons, which degrades its optical response. Both defects can couple to moderate-$Q$ subwavelength nanoscale photonic crystal cavities, fabricated in diamond itself.
An example quantum network architecture was recently realised by Knaut et al. [43], with two nodes built around SiV in diamond photonic crystal cavities, operating below 200 mK. Memory qubits were realised using nuclear spin of a nearby 29Si (with 10 ms decoherence times, controlled by RF electronics), coupled to the fiber modes via a communications qubit (electronic spin states of SiV) and a cavity with moderate quality factor $Q\sim 10^{3}$ and cooperativity $C\sim 10$. Even accounting for the additional inefficiencies, including cavity-fiber coupling, correcting the frequency mismatching between the SiVs in the two nodes, downconverting optical emission from the SiV at 737 nm closer to the telecom wavelength at 1350 nm, and losses suffered by the photonic qubits over 45 km propagation between the nodes, the this simple network realised entanglement between the two memories with fidelity of 0.7 at 1 Hz rate.
@@ V.3.2 Neutral atoms and ions
While trapped atoms are arguably the most should mature technology for realising quantum nodes, the enormous footprint of ultra-high vacuum and optical setups significantly limits their scalability, and opportunities for deployment in large-scale quantum networks.
V.4 Light-matter coupling gates
This light-matter described in the previous section must preserve the photon’s temporal and phase information, as these encode the interferometric signal of interest. Moreover, because each photon may only occupy a narrow time bin and arrive probabilistically, the entire absorption-and-storage operation must be repeatable and low-latency, with errors significantly below the mean photon number $\epsilon$ per bin. Any imperfection at this stage reduces the signal-to-noise ratio of the subsequent quantum measurement, and hence the overall sensitivity of the VLBI array.
The state-of-the-art gate fidelity and speeds for various platforms include:
  • ion traps: Single-qubit gates, error $1.5\times 10^{-6}$ per Clifford and gate time 600 ns [50], and Multi-qubit gates: Two-qubit gates at 270 $\mu s$ [51].
  • neutral-atoms : two-qubit gates via Rydberg blockade with error $5\times 10^{-3}$ and gate time $250$ns [52]; Ref. [53] reports a two-qubit logic gate using Förster resonances with error $0.01$ and gate time $6.5$ns. Single-site addressable measurements have been performed with an error of $7.3\times 10^{-3}$ over a time $36$ms [54].
  • Superconducting qubits (Google) [55]: $\sim$ 20 ns for single qubits and $\sim$ 40 ns for two-qubit gates.
The problem of storing and processing quantum information converted from optical photons into matter qubits for the purposes of quantum sensing does not require full fault tolerant quantum computing. This is because the logical operations are Clifford operations in which case it suffices to have elementary gate errors below the threshold for fault tolerant quantum error correction.
@@ V.5 Entanglement distribution and repeaters
In quantum-enabled VLBI, pre-shared entanglement replaces the direct optical link between telescope sites, enabling nonlocal measurements without transmitting fragile stellar photons across long baselines. This requires distributing high-fidelity Bell pairs between nodes at rates compatible with the time-bin structure and detection probability. The quality and rate of entanglement distribution set key limits on the overall system performance.
Some state-of-the-art entanglement generation specifications include:
  • Experimental entanglement generation: by bulk-optics and polarisation-entangled photons, achieved key rate greater than 1 Gbit/sec at telecom [56]
  • In an integrated SiC platform, entangled photon pair generation rates $\sim 10^{4}$ pairs/s [57].
  • Commercial Entanglement Sources: Commercial products, such as those from Quantum Computing Inc., offer entangled photon sources with generation rates of 5M pairs/$\mu$W/s.
  • Photonic Crystal Cavities: Utilising photonic crystal cavities, time-energy entangled photon pairs have been generated at on-chip rates up to 22 MHz using only 36 $\mu$W of pump power.  [58]
In terms of repeaters, the latest progress includes:
@@
  • Experimental demonstration of heralded entanglement between two absorptive quantum memories separated by 3.5 meters. This setup, utilizing polarisation-entangled photon-pair sources and solid-state quantum memories with up to 1 GHz bandwidth, successfully distributed maximally entangled states with a fidelity of 80.4 [61].
  • Towards a frequency-multiplexed repeater [62]: spectrally Multiplexed Photon-Pair Source; signal at 795 nm and idler at 1532 nm (telecom): up to 200 spectral modes, each mode $\sim$ 1.5 GHz (at 1532 nm), with 6.5 GHz spacing.
  • Proposal: Deterministic Quantum Repeater with Single Atoms in Cavities
Over the past decade, entanglement distribution rates have improved dramatically, driven by breakthroughs in source quality, detection efficiency, and the emergence of quantum memories and spectral/temporal multiplexing techniques. This marks an orders-of-magnitude increase in capability, especially when comparing today’s multiplexed and memory-integrated systems to earlier linear optics-only approaches.
These developments represent real progress. Demonstrations like the memory-enhanced Bell-state measurements using solid-state spins in nanophotonic cavities, or the use of spectrally multiplexed SPDC sources with quantum-memory-compatible bandwidths, have shown that entanglement generation rates can be significantly boosted, even under realistic loss conditions.
That said, we’re still far from practical, high-rate quantum repeater networks.
  • Rates remain low when memory is involved
  • Multiplexing is not fully integrated yet—many setups demonstrate multiplexed sources or filters, but don’t yet combine this with active feed-forward control and working quantum memories.
  • Most systems often require frequency conversion, which adds loss and complexity.
  • Scalability is an open challenge—interfacing multiple quantum memories, handling timing jitter, and maintaining fidelity across many links will require engineering solutions beyond current lab-scale demonstrations.
Below, we summarise the various protocols in the current literature.
@@ VI Literature survey
Several protocols have made advances to quantum-enabled VLBI, we summarise a selection of the following protocols, which are listed in approximate chronological order of appearance.
@@
  • Longer-Baseline Telescopes Using Quantum Repeaters [2].
  • Quantum memories and the double-slit experiment: implications for astronomical interferometry [63].
  • Experiment demonstration [64] of the protocol in Ref. [2]
  • Optical Interferometry with Quantum Networks: scheme based on efficient quantum data compression offers an exponential decrease in required entanglement bandwidth [3].
  • Quantum-assisted telescope arrays: frequency and multi-domain extension to [3], GHZ states used for parity checks [4]
  • Using STIRP to absorb the stellar photons’ information into the quantum memories, and incorporate quantum error correction [5].
  • Optimal qubit circuits for quantum-enhanced telescopes [65]: encodes the time of arrival of the stellar photon into a quantum memory; reduces both the number of ancilla qubits and the number of gate operations by a factor of two.
  • Vacuum beam guide for large-scale quantum networks: remove most of the quantum tech and put a variable delay line back in [6].
  • Continuous variable teleportation schemes based on two-mode squeezed-vacuum [66, 67]
  • Comparison between amplitude and intensity interferometry as a function of baseline and mean photon number [68]
  • One experimental demonstration of intensity interferometry [69]
@@
Note that $||\bm{\mathrm{I}}^{M}||$ can be significantly larger than $\epsilon$ for $\gamma\rightarrow 1$, and that care should be taken when discussing the sum of the two terms: this is because per photon, $J_{\phi}$ is upper bounded by $1$, whereas $J_{\gamma}$ can, in principle, approach infinity.
@@ VI.2 Longer-Baseline Telescopes Using Quantum Repeaters [2]
Suppose we can distribute the following entangled state,
@@
Now, each half shared state |$\ket{\psi_{\mathrm{shared}}}$ is then interfered with the stellar state at the respective telescope station with a 50:50 beam splitter (on each side, there are two detectors). The coincidence count can be used to directly directly to measure the $\gamma e^{i\phi}$. We post-select the measurement results, considering the events where a single photon is observed at $A$ and $B$ simultaneously. The total probability of seeing correlated photons at the output, conditioned on having one click at each telescope, is,
$\frac{1}{2}(1+\mathrm{Re}[\gamma e^{-i(\phi-\delta)}]),$ (27)
and the total probability of seeing an anti-correlation is,
@@
This protocol requires that a “ground” photon interferes with a stellar photon, which means that the two need to be matched in their respective temporal and frequency modes. This implies that per photon measured, the entanglement consumption is of order $1/\epsilon$: see unary encoding below.
Ref. [71] extended the analysis to distributing more ground photons, where parallel-distributed ground photons interfere with the astronomical photon in a balanced multimode beam splitter. Ref. [72] investigated the detrimental effect of having distinguishable ground photons.
@@ VI.3 Unary encoding
Let us now introduce quantum memory. Consider the simplest unary encoding where there is one memory qubit for each time bin. Given $M$ bins, after a single photon is stored, in the 5th time bin, for example, the memories will be in the state
@@ VI.4 Unary encoding with binary search
Figure 6: Schematic for binary search for the time bin into which the photon arrived. This protocol is a new result.
We can use an encoding that is efficient in the entanglement consumption, where the encoding is akin to a binary search. We choose a block size $M$ where $M\epsilon\sim 1$.
  • For each run of the parity check, we divide the memory block into two (Fig. 6).
  • To search for where the photon is within the time bin, we apply CZ gates between all of Alice’s or Bob’s memory qubits.
In our example with 4 time bins, a photon arrives in the second bin and is shared non-locally between Alice and Bob. The state is
$\rho\propto$ |$\ket{0,0}_{A_{1}B_{1}}(\ket{01}_{A_{2}B_{2}}\pm e^{i\phi}\ket{10}_{A_{2}B_{2}})$
|$\ket{0,0}_{A_{3}B_{3}}\ket{0,0}_{A_{4}B_{4}}$ (33)
We divide the block of 4 into two halves. The first half is
$\rho_{1,2}\propto$ |$\ket{0,0}_{A_{1}B_{1}}(\ket{01}_{A_{2}B_{2}}\pm e^{i\phi}\ket{10}_{A_{2}B_{2}})$ (34)
@@
$\rho_{1,2}\otimes\ket{\Phi^{+}}$ $\stackrel{{\scriptstyle 4\times\text{CZ}}}{{\rightarrow}}\rho_{1,2}\times\ket{\Phi^{-}}$ (35)
@@
$\rho_{3,4}\otimes\ket{\Phi^{+}}$ $\stackrel{{\scriptstyle 4\times\text{CZ}}}{{\rightarrow}}\rho_{3,4}\times\ket{\Phi^{+}}$ (36)
The procedure continues by iteratively dividing the block containing the photon into halves, until a single time bin remains.
This encoding is still linear in the memory requirements, where the number of memory qubits scales as $1/\epsilon$. However, the entanglement consumption has been reduced to scaling with $\sim~{}\log_{2}(1/\epsilon)$

### v2

@@ V Technological requirements of Optical VLBI
    Starlight is collected and time-multiplexed into different bins (see Fig. 4)
@@
    The collected starlight is absorbed into quantum memories. Non-local measurement is performed between quantum memories of the same time bin between the telescope sites. Per time bin, we need at least
@@
      One set of logical single-qubit gates to perform the measurements. This one is non-Clifford in general.
Given that each shared time-bin contains a state of mean photon number $\epsilon$, the logical gate error rate needs to be much smaller than $\epsilon$. Below, we outline these core requirements and the current state of the art.
@@ V.1 Phase stabilisation
Maintaining phase coherence across distant telescope sites is critical for interferometry. The entire system must be stabilised to within a small fraction of the optical wavelength. A shared optical phase reference is necessary. Notably, recent twin-field QKD experiments have achieved 98% interference visibility over 200 km at 1550 nm [37], suggesting that phase stabilisation over VLBI-scale baselines is technically feasible. A quantitative analysis on the shared phase reference is given in Ref. [38].
@@ V.2 Time-bin multiplexing
Starlight must be parsed into discrete time bins, with each bin corresponding to the coherence time of the incoming thermal light (see Fig 4), because signal separated by longer than the coherence time will not interfere . The coherence time of the light is inversely proportional to the bandwidth $\Delta\nu$ [29],
@@
The bandwidth of the stellar source is often quoted in either frequency or wavelength. As the bandwidth increases, the coherence time—and therefore the required time-bin size—decreases proportionally. As an example, for light generated by thermal sources with a narrow spectral width ($\Delta\nu\sim 10^{8}$ Hz), the corresponding coherence time is of order $10^{-8}s$, and coherence length is of order 3 m.
Current state-of-the-art optical switching operates at rates of up to 38 MHz with approximately 80% transmission efficiency [39], corresponding to a minimum time-bin size of roughly $2.6\times 10^{-8}$ seconds. In contrast, at a wavelength of 1550 nm, a spectral bandwidth of 1 nm corresponds to a temporal coherence time of about 8 picoseconds—nearly 3000 times shorter than the achievable switching interval.
@@ V.3 Implementing quantum nodes with optical cavities
The many platforms for implementing quantum nodes in physical systems are built around memory qubits, which store the information in long-lived nuclear spin, dark or metastable states, perfectly isolated from the environment. To control and read out that memory, quantum nodes would typically use a direct, controlled coupling to a cavity, or mediate it via a short-lived communication qubit [41]. The cavity is used to boost the intrinsically low efficiency of these processes and enhance the in and out-coupling from the memory to the fibre.
To this end, the cavity operates in the over coupled regime, meaning that its coupling to the fibre mode $\kappa_{\text{out}}$ should dominate over other losses $\kappa_{0}$, and support high cooperativity $C=g^{2}/\kappa\gamma\gg 1$, defined by the cavity-qubit coupling $g$, and decoherence rate of the qubit $\gamma$.
@@ V.3.1 Atomic defects, incorporating nitrogen, silicon, and germanium atoms
Atomic defects, incorporating nitrogen, silicon, or germanium atoms next to vacancies in diamond (NV, SiV, and GeV) can implement both the communication and memory qubits, encoding information in the electronic transitions of the defect, nuclear spins of the nearby atoms—13C for NVs, and 29Si for SiVs—respectively [43, 44]. These two types of defects also offer very different challenges and opportunities: NV exhibits far longer coherence times at room temperature ($\sim 1$ s, several orders of magnitude over SiV), but also is far more sensitive to electrical noise and coupling to phonons, which degrades its optical response. Both defects can couple to moderate-$Q$ subwavelength nanoscale photonic crystal cavities, fabricated in diamond itself.
An example quantum network architecture was recently realised by Knaut et al. [45], with two nodes built around SiV in diamond photonic crystal cavities, operating below 200 mK. Memory qubits were realised using the nuclear spin of a nearby 29Si (with 10 ms decoherence times, controlled by RF electronics), coupled to the fiber modes via a communications qubit (electronic spin states of SiV) and a cavity with moderate quality factor $Q\sim 10^{3}$ and cooperativity $C\sim 10$. Even accounting for the additional inefficiencies, including cavity-fibre coupling, correcting the frequency mismatching between the SiVs in the two nodes, downconverting optical emission from the SiV at 737 nm closer to the telecom wavelength at 1350 nm, and losses suffered by the photonic qubits over 45 km propagation between the nodes, the this simple network realised entanglement between the two memories with fidelity of 0.7 at 1 Hz rate.
@@ V.3.2 Neutral atoms and ions
While trapped atoms are arguably the most mature technology for realising quantum nodes, the enormous footprint of ultra-high vacuum and optical setups significantly limits their scalability, and opportunities for deployment in large-scale quantum networks.
V.4 Quantum gates
This light matter described in the previous section must preserve the photon’s temporal and phase information, as these encode the interferometric signal of interest. Moreover, because each photon may only occupy a narrow time bin and arrive probabilistically, the entire absorption-and-storage operation must be repeatable and low-latency, with errors significantly below the mean photon number $\epsilon$ per bin. Any imperfection at this stage reduces the signal-to-noise ratio of the subsequent quantum measurement, and hence the overall sensitivity of the VLBI array.
Recent experimental advances have demonstrated high gate fidelity and speeds for various platforms include
  • Ion traps: Single-qubit gates with error $1.5\times 10^{-6}$ and 600 ns gate time [52]; two-qubit gates in 270 $\mu$[53].
  • Neutral atoms: Rydberg-blockade gates with $\sim 5\times 10^{-3}$ error in 250 ns [54]; Förster resonance gates achieving 1% error in 6.5 ns [55]; single-site readout errors $\sim 7\times 10^{-3}$ over 36 ms [56].
  • Superconducting qubits: Single- and two-qubit gate times of 20 ns and 40 ns, respectively, as demonstrated by Google [57].
Importantly, quantum sensing in this context does not require universal, fault-tolerant quantum computation. Maintaining gate errors below standard quantum error correction thresholds is sufficient for reliable operation.
@@ V.5 Entanglement distribution and repeaters
In quantum-enabled VLBI, pre-shared entanglement substitutes for direct optical links between telescope sites, enabling nonlocal measurements without transmitting fragile stellar photons across long baselines. The performance of such a system is fundamentally limited by the rate and fidelity of Bell pair distribution.
Recent progress in entangled photon sources includes polarization-entangled photons at telecom wavelengths have enabled key rates exceeding 1 Gbit/s [58]; in SiC systems, pair generation rates of $\sim 10^{4}$ pairs/s have been demonstrated[59]; commercial sources: off-the-shelf systems, such as those from Quantum Computing Inc., now reach 5M pairs/$\mu$W/s; on-chip generation of time-energy entangled pairs has achieved 22 MHz using only 36 $\mu$W of pump power  [60].
@@
These developments mark a substantial progress from earlier linear-optics-only approaches, thanks to improvements in source quality, detection efficiency, and the emergence of quantum memories and spectral/temporal multiplexing techniques. However, significant challenges remain before practical, high-rate repeater networks are feasible:: rates remain low when memory is involved; multiplexing is not fully integrated yet-—many setups demonstrate multiplexed sources or filters, but do not combine this with active feed-forward control and working quantum memories; most systems often require frequency conversion, which adds loss and complexity; scalability is an open challenge—interfacing multiple quantum memories, handling timing jitter, and maintaining fidelity across many links will require engineering solutions beyond current lab-scale demonstrations.
@@ VI Literature survey
Many protocols have made advances to quantum-enabled VLBI, we summarise a selection of the following; these are listed in approximate chronological order.
@@
  • Longer-Baseline Telescopes Using Quantum Repeaters [2], and an experimental demonstration [66].
  • Optical Interferometry with Quantum Networks [3].
  • Quantum-assisted telescope arrays [4]: frequency and time-multiplexed extension to [3].
  • Quantum memories and the double-slit experiment (hard drive-based): implications for astronomical interferometry [67].
  • Incorporating quantum error correction [5].
  • Optimal qubit circuits for quantum-enhanced telescopes [68].
  • Vacuum beam guide for large-scale quantum networks [6].
  • Continuous variable teleportation schemes based on two-mode squeezed-vacuum [36, 69]; and fundamental limits in the absence of non-Gaussian resources [70].
  • Comparison between amplitude and intensity interferometry as a function of baseline and mean photon number [71], see also the comparison with “weak heterodyne scheme” [72].
  • Experimental demonstration of intensity interferometry [73]
@@
Note that $||\bm{\mathrm{I}}^{M}||$ can be significantly larger than $\epsilon$ for $\gamma\rightarrow 1$.
@@ VI.2 Longer-Baseline Telescopes Using Quantum Repeaters [2]
Suppose we can distribute the following entangled state (“ground” photon),
@@
Now, each half of |$\ket{\psi_{\mathrm{shared}}}$ is then interfered with the stellar state in Eq. (IV) at the respective telescope station with a 50:50 beam splitter. Then, we post-select and consider the events where a two-photon coincidence is observed at stations $A$ and $B$ simultaneously. The total probabilities of seeing photons in correlated, and anticorrelated outputs are respectively
$\frac{1}{2}(1+\mathrm{Re}[\gamma e^{-i(\phi-\delta)}]),$
@@
This protocol requires the “ground” photon to to interfere with a stellar photon, which means that the two need to be matched in their respective temporal and frequency modes. This implies that per stellar photon measured, the entanglement consumption is of order $1/\epsilon$: see unary encoding below.
Ref. [74] discusses a multi-photon extension of Ref. [2]. This is complementary to Ref. [75], which extended the analysis to distributing more ground photons, where parallel-distributed ground photons interfere with the astronomical photon in a balanced multimode beam splitter. Ref. [76] investigated the detrimental effect of having distinguishable ground photons.
@@ VI.3 Unary encoding
Let us now introduce quantum memory. Consider the simplest unary encoding where there is one memory qubit for each time bin. Suppose there are $M$ time bins in total. After a single photon is stored, in the 5th time bin, for example, the memories will be in the state
@@ VI.4 Unary encoding with binary search
Figure 6: Schematic for binary search for the time bin into which the photon arrived.
We can use an encoding that is efficient in the entanglement consumption, where the encoding is akin to a binary search. We can choose a block size $M$ such that $M\epsilon\sim 1$. For each run of the parity check, we divide the memory block into two (Fig. 6). To search for where the photon is within the time bin, we apply CZ gates between all of Alice’s or Bob’s memory qubits.
In our example with 4 bins, a photon arrives in the second bin and is shared non-locally between Alice and Bob. The state is
$\rho=$ |$\ket{0,0}\bra{0,0}_{A_{1}B_{1}}\otimes\rho^{\star^{\prime}}_{A_{2}B_{2}}\otimes$
[…]
$\rho^{\star^{\prime}}$ $=\left(\frac{1+\gamma}{2}\right)\ket{\psi_{+}^{\phi}}\bra{\psi_{+}^{\phi}}+\left(\frac{1-\gamma}{2}\right)\ket{\psi_{-}^{\phi}}\bra{\psi_{-}^{\phi}}$ | (32)
We divide the block of 4 time bins into two halves. The first half is
$\rho_{1,2}=$ |$\ket{0,0}\bra{0,0}_{A_{1}B_{1}}\otimes\rho^{\star^{\prime}}_{A_{2}B_{2}}$ (33)
@@
$\rho_{1,2}\otimes\ket{\Phi^{+}}\bra{\Phi^{+}}$ | $\stackrel{{\scriptstyle 4\times\text{CZ}}}{{\rightarrow}}\rho_{1,2}\otimes\ket{\Phi^{-}}\bra{\Phi^{-}}$ | (34)
@@
$\rho_{3,4}\otimes\ket{\Phi^{+}}\bra{\Phi^{+}}$ | $\stackrel{{\scriptstyle 4\times\text{CZ}}}{{\rightarrow}}\rho_{3,4}\otimes\ket{\Phi^{+}}\bra{\Phi^{+}}$ | (35)
The procedure continues by iteratively dividing the block containing the photon into halves, until a single time bin remains. This encoding is linear in the memory requirements, where the number of memory qubits scales as $1/\epsilon$. However, the entanglement consumption has been reduced to $\sim\log_{2}(1/\epsilon)$.